# Transition to Petschek Reconnection in Subrelativistic Pair Plasmas: Implications for Particle Acceleration

### v1

@@ 3.2 Single to Multiple X-point Transition
To see why the $\sigma=10$ case is plasmoid-dominated and the $\sigma=0.3$ case is exhaust-dominated, let us consider spacetime diagrams showing the birth, growth, and outflow of plasmoids in the current sheet (Figure 6). The key difference between the two regimes is in plasmoid mergers, the process by which two plasmoids collide and form a larger plasmoid (for extensive analysis, see L. Sironi et al., 2016). In the $\sigma=10$ case (Figure 6(a)) this occurs constantly, with each plasmoid undergoing many mergers within its lifetime. In contrast, in the $\sigma=0.3$ case (Figure 6(b)) plasmoids seldom collide – the vast majority of plasmoids are simply born and then advect out of the diffusion region. Without mergers, these plasmoids never become very large – they simply grow modestly and then leave at the Alfvén speed (see also Figure 2).
@@
Figure 6: Spacetime plots of the midline ($x=L_{x}/2=60\;c/\omega_{p}$) for Runs I and II respectively after they have reached equilibrium. The areas with the lowest density (the darkest blue) are the diffusion regions and the high-density streaks of yellow are the plasmoids – thus these plots show the life journey of plasmoids as they are born and leave the domain. As one can see in the $\sigma=0.3$ case, plasmoid production, while regular, is sufficiently infrequent to allow most plasmoids to escape the region near the X-point at the center without ever merging with another plasmoid (though one prominent collision can be seen near (190, 180)). The case in Run I is quite different - here we see rapid plasmoid production and, crucially, many mergers. This results in large, slow plasmoids which themselves are more susceptible to further mergers. This feedback mechanism fills the domain with large, massive plasmoids.
While the steady stream of numerous large plasmoids in the $\sigma=10$ case swiftly disrupts any nascent exhausts, the trickle of small plasmoids in the $\sigma=0.3$ is sufficiently feeble that exhausts can form unimpeded. Although the plasmoids in the latter case do perturb the exhaust as they pass through (see Figure 1b and M. E. Innocenti et al. (2015)), they are too few in number to permanently disperse it. Furthermore, these effects are self-reinforcing. The plasmoid-dominated current sheet that forms in the high-$\sigma$ case spawns secondary X-points which produce yet more plasmoids. The wide-angle exhausts which form in the low-$\sigma$ case thicken the current sheet away from the X-point, precluding further tearing. The plasmoids passing through the exhausts, immersed in steady flow, neither accrete fresh flux nor merge with each other. Instead, they are simply swept out of the system.
@@
To quantify this phase transition, we define a dimensionless parameter $\mathcal{P}$ which scales as the ratio between the plasmoid width and the inter-plasmoid distance, which we term the plasmoid parameter. As $\mathcal{P}$ rises, so too does the chance for mergers. Hence for high $\mathcal{P}$ we expect multiple X-point reconnection, and for low $\mathcal{P}$ we expect instead single X-point reconnection. If we define $\mathcal{F}$ as the plasmoid production rate at the X-point measured in plasmoids per unit time, the mean inter-plasmoid distance will scale as $v_{A}/\mathcal{F}$. The width of a plasmoid upon leaving the diffusion region is roughly that of the current sheet itself, $w_{cs}$. Thus our plasmoid parameter may be expressed as
@@
The approximate parameters as determined from Runs I and II and Equation (10) are given in Table 2. We can conclude that $\mathcal{P}>0.10$ corresponds to multiple X-point reconnection and $\mathcal{P}<0.05$ to single X-point, separated by a critical transition parameter $\mathcal{P}_{c}\sim 0.075$. As $\mathcal{P}$ increases from $0.05$ to $0.10$ the stream of plasmoids intensifies as more mergers occur until any potential exhaust is rapidly dispersed. Since this expression is not unique to the particular system properties, we expect it to hold in electron-ion plasmas as well. Similar current sheet structures are indeed observed in such cases (e.g., M. E. Rowan et al., 2017).
Run $\sigma$ $\mathcal{F}$ $\omega_{p}$] $w_{cs}$ $c/\omega_{p}$] $v_{A}$ [$c$] $\mathcal{P}$
I 10 0.08 1.3 0.95 0.10
II 0.3 0.02 1.4 0.48 0.05
Table 2: Key plasmoid parameters (Runs I and II). Note that we calculate $\mathcal{F}$ as the rate at which fully-developed plasmoids leave the diffusion region per unit time.
@@ 4 Particle Acceleration
Our motivation for investigating the transition from the multiple X-point regime to the single X-point regime is to assess the effect of changes in the current sheet geometry on particle acceleration mechanisms. Now we discuss the particle energy spectrum in both cases and explain why they differ – in particular, how the energization pathways in the $\sigma=0.3$ case differ from those in the canonical $\sigma=10$ case.
@@ 4.1 Total Spectra
The electron energy spectra – specifically, the electron kinetic energy normalized to the magnetization – are compared in Figure 7. In the $\sigma=10$ case (e.g., L. Sironi & A. Spitkovsky, 2014; F. Guo et al., 2014), the plasmoid-dominated reconnection results in the establishment of a broad, hard power law (that is, at energies greater than those of the cold inflow component seen on the left of Figure 7). For the $\sigma=0.3$ case, we see two distinct components in the energized part corresponding to particles within the current sheet. First, on the left, we see a shoulder comprised of a flat spectrum with a shoulder. At the end we see a small high-energy tail containing the most energetic particles which branches off from the shoulder at $\gamma\approx 1.7$. These features are caused by distinct particle acceleration mechanisms: the first “shoulder” component consists of particles that are energized in the exhaust and the second more energetic component consists of particles confined in plasmoids.
@@
Figure 7: Electron energy spectra for Runs III and IV for the entire domain, normalized by the magnetization. The transition between the exhaust component and the plasmoid component is indicated by the dotted blue line. Note that the first structures on the far left for each case represent the inflow which has acquired a small thermal temperature due to wave heating. As the domain size ($L_{y}$) grows for $\sigma=0.3$, the shape of the plot looks the same however the magnitude of the plasmoid component rapidly decreases from what is shown here due to the small fraction of the domain occupied by plasmoids (c.f. Fig 3).
@@
In relativistic reconnection, particle acceleration is fundamentally stochastic – particles can transition from direct X-point acceleration to Fermi acceleration within plasmoids and back again. This results in the development of a smooth power-law spectrum. By contrast, in the $\sigma=0.3$ case, the regular, laminar structure of the reconnection region results in two distinct acceleration pathways which depend on where the particle first encounters the current sheet. The first category of particles follows the field lines into the diffusion region. At the X-point they undergo Speiser motion (T. W. Speiser, 1965) and are accelerated by the non-ideal reconnection electric field. They are then swiftly confined by a growing plasmoid. These particles then continue to gain energy via the ideal electric fields therein as the plasmoid accretes flux and compresses (H. Hakobyan et al., 2021). These particles constitute the most energetic part of the final spectrum, the exponential cutoff with $\gamma>1.7$. This will be discussed in more detail in Section 4.2. The much larger second category of particles, instead of first encountering the current sheet at the X-point, do so at the slow-mode shocks at the boundary of the exhaust outflow. There they are pickup-accelerated and brought up to speed with the Alfvénic outflow, resulting in a thermal distribution. This will be discussed in Section 4.3. Since the X-point occupies a very small portion of the overall domain, the vast majority of particles are of the second kind, and they absorb nearly all of the liberated magnetic free energy. This dominance only grows as one considers realistic system sizes where $L_{y}\gg c/\omega_{p}$.
@@ 4.2 X-point Injected Particles
Recent work on particle acceleration in relativistic reconnection has demonstrated that the most energetic particles undergo a two-step acceleration process: injection to relativistic energies via direct acceleration at X-points and then strong energization by Fermi/betatron acceleration powered by ideal electric fields in plasmoids (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025). Some particles undergo an analogous process in $\sigma=0.3$ reconnection. Particles which fall into the X-point are accelerated by the non-ideal electric field as they undergo Speiser motion. The result is the establishment of a power-law distribution in the vicinity of the X-point. These particles then are quickly trapped by plasmoids where they ultimately gain most of their energy by Fermi/betatron acceleration as the plasmoid compresses and its magnetic fields intensify.
Let us first consider the process by which the initial power law is formed at the X-point for $\sigma=0.3$ and, specifically, how it differs from that of the $\sigma=10$ case. We follow the procedure of N. Bessho & A. Bhattacharjee (2012) (hereafter BB12) who derived the exit energy spectrum of particles leaving the X-point in the fully-relativistic case. Since the amount of energy gained in the diffusion region depends on the total amount of time the particle remains undergoing Speiser motion before being advected out by $B_{x}$, particles which enter closer to the center where $B_{x}$ is weaker will remain longer and hence attain higher energies than those which enter closer to the edge of the diffusion region.
@@
To ensure a fair comparison, we compare the primary X-point in the $\sigma=0.3$ case (here the only X-point) with the primary X-point in the $\sigma=10$ case. In BB12 the authors define the edge of the diffusion region as the $E=B$ contour, such that the diffusion region is where the electric field dominates over the magnetic field. This is unsuitable in the $\sigma=0.3$ case since the $E>B$ zone only covers a small sliver of the diffusion region. We use the modified boundary $E=(v_{A}/c)B$ since this will define the region in which the electric field term in the Lorentz force law will be greater than that of the magnetic field for a particle leaving the diffusion region at the Alfvén speed. Since for $\sigma=10$ the Alfvén speed is highly relativistic, $v_{A}=0.95c$, this leads to nearly the same condition for the high-magnetization case but correctly allows for a larger diffusion region for low magnetization (Figure 8). As in BB12, we consider a particle as having exited the diffusion region if it passes the maximum or minimum $y$-value of the $E=(v_{A}/c)B$ contour at that point in time. The exit spectrum is thus the time-integrated spectrum of particle energies as they cross this boundary.
@@
Figure 8: X-points from Runs I and II. The non-ideal electric field $\mathbf{E}_{ni}\equiv\mathbf{E}+\mathbf{v}\times\mathbf{B}/c$ in the out-of-plane direction is plotted. The $E=(v_{A}/c)B$ contours, giving the boundaries of the diffusion region for the purposes of computing the exit spectra, are shown in yellow.
@@
The exit spectra are shown in Figure 9. As one can see, in both cases we observe hard spectra – there does not seem to be any diminished ability by the subrelativistic diffusion region to generate a power-law spectrum. In fact, particles gain slightly more energy in the $\sigma=0.3$ case relative to the magnetic free energy available. This supports our presumption that there is nothing inherently special about $E>B$ regions for the non-relativistic case, since particles can still effectively sample the non-ideal field while $B>E$, since $v/c$ is correspondingly lower.
@@
Figure 9: Normalized kinetic energy values for electrons upon their departure from the primary diffusion region, summed over time. The dotted line shows the analytical expression for the nonrelativistic exit spectrum (Equation (14)). The humps seen in the curve for $\sigma=0.3$ are caused by a Hall-like $E_{x}$ component noted by M. Swisdak et al. (2008) which results in step-wise acceleration and bunches up the electron distribution; see Appendix B for details. From left to right, the peaks correspond to particles which complete one, two, and three Speiser orbits before leaving the diffusion region.
@@
If we, like BB12, neglect the Speiser motion in the $x$-direction and repeat their procedure to compute an analytical form for the exit spectrum in the non-relativistic limit, we arrive at the following result which resembles that of S. V. Bulanov & P. V. Sasorov (1976):
@@
Here $\varepsilon\equiv\sqrt{qE_{z}^{2}/bmc^{2}}$ is a dimensionless parameter quantifying the relative influence of $E_{z}$ on the particle trajectory compared to $B_{y}$, $W^{*}\equiv(qE_{z}^{2}/b)^{-1}W$ is the normalized kinetic energy, and $b\equiv\partial B_{x}/\partial y$, assumed to be constant. Details of the derivation can be found in Appendix A. This fits the data neatly (Figure 9) and may further explain the experimental observations of exponential cutoffs in electron energy spectra (W. Fox et al., 2010; B. Na et al., 2023).
Figure 10: Trajectory of an electron over the density profile for the last timestep plotted for Run I ($\sigma=0.3$). $\Omega$ is the upstream electron gyrofrequency. The particle first interacts with the current sheet at the X-point and is thereafter confined within a plasmoid where the initial direct $E$-field energization is magnified several times by Fermi/betatron heating. The oscillation in energy reflects the gyration up and down the $E_{z}$ potential.
Once this power law is established, it is enhanced by acceleration within plasmoids. Consider the particle shown in Figure 10, a typical case. This electron gains its first substantial amount of energy upon interaction with the X-point at $\Omega t\approx 100$. However, most of its ultimate energy gain comes from being confined within a plasmoid as it compresses and its magnetic fields intensify (F. Guo et al., 2019; H. Hakobyan et al., 2021). This energization can persist so long as the plasmoid continues to accrete more flux and compress adiabatically, heating the particles within. Once the plasmoids are completely immersed within the exhaust, they are deprived of fresh magnetic flux, starving this acceleration mechanism of additional free energy. As a result, the vast majority of plasmoid heating occurs while the plasmoid is still close to the X-point: past $\gtrsim 1000\;c/\omega_{p}$ skin depths the plasmoids no longer grow and the particles within the plasmoid subsequently no longer gain energy. Thus, unlike in relativistic reconnection, this acceleration pathway has a fixed window of opportunity and cannot generate arbitrarily high-energy particles provided a large enough domain.

### v2

@@ 3.2 Single to Multiple X-point Transition
To see why the $\sigma=10$ case is plasmoid-dominated and the $\sigma=0.3$ case is exhaust-dominated, let us consider spacetime diagrams showing the birth, growth, and outflow of plasmoids in the current sheet (Figure 4). The key difference between the two regimes is in plasmoid mergers, the process by which two plasmoids collide and form a larger plasmoid (for extensive analysis, see L. Sironi et al., 2016). In the $\sigma=10$ case (Figure 4(a)) this occurs constantly, with each plasmoid undergoing many mergers within its lifetime. In contrast, in the $\sigma=0.3$ case (Figure 4(b)) plasmoids seldom collide—the vast majority of plasmoids are simply born and then advect out of the diffusion region. Without mergers, these plasmoids never become very large—they simply grow modestly and then leave at the Alfvén speed (see also Figure 2). While the steady stream of numerous large plasmoids in the $\sigma=10$ case swiftly disrupts any nascent exhausts, the trickle of small plasmoids in the $\sigma=0.3$ is sufficiently feeble that exhausts can form unimpeded.
@@
Figure 4: Spacetime plots of the midline ($x=L_{x}/2=60\;c/\omega_{p}$) for Runs I and II respectively after they have reached equilibrium. The areas with the lowest density (the darkest blue) are the diffusion regions and the high-density streaks of yellow are the plasmoids—thus these plots show the life journey of plasmoids as they are born and leave the domain. As one can see in the $\sigma=0.3$ case, plasmoid production, while regular, is sufficiently infrequent to allow most plasmoids to escape the the X-point at the center without ever merging with another plasmoid (though one prominent collision can be seen near (190, 180)). The case in Run I is quite different - here we see rapid plasmoid production and, crucially, many mergers. This results in large, slow plasmoids which themselves are more susceptible to further mergers. This feedback mechanism fills the domain with large, massive plasmoids.
@@
To quantify this phase transition, we define a dimensionless parameter $\mathcal{P}$ which scales as the ratio between the plasmoid width and the inter-plasmoid distance, which we term the plasmoid parameter. As $\mathcal{P}$ rises, so too does the chance for mergers. Hence for high $\mathcal{P}$ we expect multiple X-point reconnection, and for low $\mathcal{P}$ we expect instead single X-point reconnection. If we define $\mathcal{F}$ as the plasmoid production rate at the X-point measured in exiting plasmoids per unit time, the mean inter-plasmoid distance will scale as $v_{A}/\mathcal{F}$. The width of a plasmoid upon leaving the diffusion region is roughly that of the current sheet itself, $w_{cs}$. Thus our plasmoid parameter may be expressed as
@@
Using the measured values of $\mathcal{F}$ and $w_{cs}$, we find $\mathcal{P}=0.10$ for Run I ($\sigma=10$) and $\mathcal{P}=0.05$ for Run II ($\sigma=0.3$). We can thus conclude that $\mathcal{P}\gtrsim 0.10$ corresponds to multiple X-point reconnection and $\mathcal{P}\lesssim 0.05$ to single X-point, separated by a critical transition parameter $\mathcal{P}_{c}\sim 0.075$. As $\mathcal{P}$ increases from $0.05$ to $0.10$ the stream of plasmoids intensifies as more mergers occur until any potential exhaust is rapidly dispersed. Since this expression is not unique to the particular system properties, we expect it to hold in electron-ion plasmas as well where similar current sheet structures are indeed observed (e.g., M. E. Rowan et al., 2017).
@@ 4 Particle Acceleration
Our motivation for investigating the transition from the multiple X-point regime to the single X-point regime is to assess the effect of changes in the current sheet geometry on particle acceleration. Now we discuss the particle energy spectrum in both cases and explain why they differ—in particular, how the energization pathways in the $\sigma=0.3$ case differ from those in the canonical $\sigma=10$ case.
@@ 4.1 Total Spectra
The electron energy spectra—specifically, the electron kinetic energy normalized to the magnetization—are compared in Figure 5. In the $\sigma=10$ case (e.g., L. Sironi & A. Spitkovsky, 2014; F. Guo et al., 2014), the plasmoid-dominated reconnection results in the establishment of a broad, hard power law (that is, at energies greater than those of the cold inflow component seen on the left of Figure 5). For the $\sigma=0.3$ case, we see two distinct components in the energized part corresponding to particles within the current sheet. First, on the left, we see a shoulder comprised of a flat spectrum followed by a rapid drop-off. At the end we see a small high-energy tail containing the most energetic particles which branches off from the shoulder at $\gamma\approx 1.7$. These features are caused by distinct particle acceleration mechanisms: the first “shoulder” component consists of particles that are energized in the exhaust and the second, more energetic, component consists of particles confined in plasmoids.
@@
Figure 5: Electron energy spectra for Runs III and IV for the entire domain, normalized by the magnetization. The dashed blue line indicates the expected pickup energy, $mv_{A}^{2}$. The transition between the exhaust component and the plasmoid component is indicated by the yellow line. Note that the first structures on the far left for each case represent the inflow which has acquired a small thermal temperature due to wave heating. As the domain size ($L_{y}$) grows for $\sigma=0.3$, the shape of the plot looks the same, however the magnitude of the plasmoid component in the spectrum decreases relative to that of the exhaust due to the decreasing fraction of the domain occupied by plasmoids.
@@
In relativistic reconnection, particle acceleration is fundamentally stochastic—particles can transition from direct X-point acceleration to Fermi acceleration within plasmoids and back again. This results in the development of a smooth power-law spectrum. By contrast, in the $\sigma=0.3$ case, the regular, laminar structure of the reconnection region results in two distinct acceleration pathways which depend on where the particle first encounters the current sheet. The first category of particles follows the field lines into the diffusion region. There they are accelerated by the non-ideal electric fields (N. Bessho & A. Bhattacharjee, 2012) before being confined and energized by a compressing plasmoid (H. Hakobyan et al., 2021). The much larger second category of particles, instead of first encountering the current sheet at the X-point, does so at the slow-mode shocks at the boundary of the exhaust outflow. There they are pickup-accelerated and brought up to speed with the Alfvénic outflow, resulting in a thermal distribution. It is this latter pathway which ultimately dominates, especially when one considers realistic system sizes where $L_{y}\gg c/\omega_{p}$.
@@ 4.2 X-point Injected Particles
Recent work on particle acceleration in relativistic reconnection has demonstrated that the most energetic particles undergo a two-step acceleration process: injection to relativistic energies via direct acceleration at X-points and then strong energization by Fermi/betatron acceleration powered by ideal electric fields in plasmoids (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025). Some particles which fall into the x-point undergo an analogous process in $\sigma=0.3$ reconnection, accelerated by the non-ideal electric field as they undergo Speiser motion (T. W. Speiser, 1965). The result is the establishment of a power-law distribution in the vicinity of the X-point. These particles then are quickly trapped by plasmoids where they ultimately gain most of their energy by Fermi/betatron acceleration as the plasmoid compresses and its magnetic fields intensify.
@@
Let us first consider the process by which the initial power law is formed at the X-point for $\sigma=0.3$ and, specifically, how it differs from that of the $\sigma=10$ case. We follow the procedure of N. Bessho & A. Bhattacharjee (2012) (hereafter BB12) who derived the exit energy spectrum of particles leaving the X-point in the fully-relativistic case. Particles which enter at the center of the diffusion region take the longest to be advected out by the $B_{x}$ field and thus gain the most energy, while those entering near the edges gain the least. To ensure a fair comparison, we compare the primary X-point in the $\sigma=0.3$ case (here the only X-point) with the primary X-point in the $\sigma=10$ case. In BB12 the authors define the edge of the diffusion region as the $E=B$ contour, such that the diffusion region is where the electric field dominates over the magnetic field. This is unsuitable in the $\sigma=0.3$ case since there the $E>B$ zone only covers a small sliver of the diffusion region. We use the modified boundary $E=(v_{A}/c)B$ since this will define the region in which the electric field term in the Lorentz force law will be greater than that of the magnetic field for a particle leaving the diffusion region at the Alfvén speed (Figure 6). As in BB12, we consider a particle as having exited the diffusion region if it passes the maximum or minimum $y$-value of the $E=(v_{A}/c)B$ contour at that point in time. The exit spectrum is thus the time-integrated distribution of exit energies.
@@
Figure 6: X-points from Runs I and $I$. The non-ideal electric field $\mathbf{E}_{ni}\equiv\mathbf{E}+\mathbf{v}\times\mathbf{B}/c$ in the out-of-plane direction is plotted. The $E=(v_{A}/c)B$ contours, giving the boundaries of the diffusion region for the purposes of computing the exit spectra, are shown in yellow.
@@
The exit spectra are shown in Figure 7. As one can see, in both cases we observe hard spectra—subrelativistic X-points accelerate particles as effectively as relativistic ones do. In fact, particles gain slightly more energy in the $\sigma=0.3$ case relative to the magnetic free energy available. This supports our presumption that there is nothing inherently special about $E>B$ regions for the nonrelativistic case, since particles can still effectively sample the non-ideal field while $B>E$, since $v/c$ is correspondingly lower.
@@
Figure 7: Normalized kinetic energy values for tracked simulation electrons upon their departure from the primary diffusion region, summed over time. The dotted line shows the analytical expression for the nonrelativistic exit spectrum (Equation (15)). The humps seen in the curve for $\sigma=0.3$ are caused by a Hall-like $E_{x}$ component noted by M. Swisdak et al. (2008) which results in step-wise acceleration and bunches up the electron distribution; see Appendix C for details. From left to right, the peaks correspond to particles which complete one, two, and three Speiser oscillations before leaving the diffusion region.
@@
If we, like BB12, neglect the Speiser motion in the $x$-direction and repeat their procedure to compute an analytical form for the exit spectrum in the nonrelativistic limit, we arrive at the following result which resembles that of S. V. Bulanov & P. V. Sasorov (1976):
@@
Here $\varepsilon\equiv\sqrt{qE_{z}^{2}/bmc^{2}}$ is a dimensionless parameter quantifying the relative influence of $E_{z}$ on the particle trajectory compared to $B_{y}$, $W^{*}\equiv(qE_{z}^{2}/b)^{-1}W$ is the normalized kinetic energy, and $b\equiv\partial B_{x}/\partial y$, assumed to be constant. Details of the derivation can be found in Appendix B. This fits the data neatly (Figure 7) and may further explain experimental observations of exponential cutoffs in electron energy spectra (W. Fox et al., 2010; B. Na et al., 2023).
Figure 8: Trajectory of an electron over the density profile for the last timestep plotted for Run I ($\sigma=0.3$). $\Omega$ is the upstream electron gyrofrequency. The particle first interacts with the current sheet at the X-point ($105\lesssim\Omega t\lesssim 120$) and is thereafter confined within a plasmoid ($\Omega t\gtrsim 120$). As the plasmoid and trapped electron move together into the outflow, the electron energy initially imparted by direct $E$-field energization is magnified several times by Fermi/betatron heating. The oscillation in energy reflects the gyration up and down the $E_{z}$ potential.
Once this power law is established, it is enhanced by acceleration within a plasmoid. Consider the particle shown in Figure 8, a typical case. This electron gains its first substantial amount of energy upon interaction with the X-point. However, most of its ultimate energy gain comes from being confined within a single plasmoid as it compresses and its magnetic fields intensify (F. Guo et al., 2019; H. Hakobyan et al., 2021). This energization can persist so long as the plasmoid continues to accrete more flux and compress adiabatically, heating the particles within. Once the plasmoid is completely immersed within the exhaust, it is deprived of fresh magnetic flux, starving this acceleration mechanism of additional free energy. As a result, the vast majority of plasmoid heating occurs while the plasmoid is still close to the X-point: past $\gtrsim 1000\;c/\omega_{p}$ skin depths, the plasmoids no longer grow and the particles within the plasmoid subsequently no longer gain energy. Thus, unlike in relativistic reconnection, this acceleration pathway has a fixed window of opportunity and cannot generate arbitrarily high-energy particles provided a large enough domain.